# Benchmarking General Purpose In-Context Learning

### v1

@@ Abstract
In-context learning (ICL) capabilities is becoming increasingly appealing towards building general intelligence. Taking this concept one step further, we draw a parallel to humans and many animals, who inherit primarily learning capabilities but refine their memory and acquire diverse skills and knowledge through extensive lifelong experiences. This parallel inspires our approach to general purpose in-context learning (GPICL). This paper introduces two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark encompasses a wide range of diverse tasks characterized by generation and interaction, minimal transferable knowledge, and long-term dependency. These features present significant challenges for models that primarily rely on context or interactions to enhance their proficiency. We hope that these benchmarks will not only advance research in GPICL but also contribute significantly to the broader field of general intelligence.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds to: fanwang.px@gmail.com and forrest@ustc.edu.cn00footnotetext: source code available at https://github.com/FutureAGI/L3C
@@ 1 Introduction
The community has long recognized a significant gap between artificial intelligence and biological intelligence, particularly regarding innate abilities and the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). For example, the initial limited capabilities of mammal infants stand in stark contrast to their skills in lifelong learning and adaptation. This disparity is further highlighted by the phenomenon known as the genomics bottleneck. Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch & Schmidhuber, 2021; Wang et al., 2022). A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 2. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
@@
Why GPICL? The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models. First, inspired by the genomics bottleneck and building on previous work (Kirsch & Schmidhuber, 2021; Wang et al., 2022), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills. Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model. In this context, GPICL could offer an ideal solution: since the meta-training stage can result in the loss of much task-specific knowledge and skills but preserves the core learning functionality, it helps prevent the meta-model from being overly influenced by non-factual information and bias in those tasks. Third, in practice, a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, making it challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
Building on those points, we may futher propose an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans. A visual illustration of this proposed pipeline with existing methodologies is depicted in Figure 2.
@@
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL.
@@ 2.3.1 Towards Meta-Language Model
We test the ability of generation of the meta-language model on two rudimentary real-world language tasks: English vocabulary learning and basic mathematical operations (addition and subtraction), which are typical for primary school education. For the English vocabulary task, the vocabulary consists of alphabet characters and common punctuation marks; for the math task, it includes digits and mathematical operators. The input contexts for testing included 40 English words and simple arithmetic operations below $5$. To discourage direct replication, we intentionally introduced errors into $15\%$ of the words or equations. We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts. The results, depicted in Figure 4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings. Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
@@ 2.3.2 Scaling Laws for GPICL
We have conducted further research into how the scale of parameters affects GPICL capability. As illustrated in Figure 5, our findings reveal notable trends: 1. Language complexity ($n$) influences the relative performance of different-sized models and affects the dependence on context length. In simpler tasks ($n<4$), performance rapidly converges to lower bounds asymptotically ($t\sim T=4k$. Here, with a slight abuse of notation, $T$ does not necessarily represent the ICL horizon, but rather denotes the maximum context length used in our experiments.), and the performances of various models become indistinguishable, despite starting differently. 2. It should be noted that for tasks as straightforward as $n=2$, where a 300K model achieves comparable asymptotic performance to a 150M model, we observe that larger models cannot attain higher performance without sufficient context. It is reasonable to infer that increasing model complexity further will not be beneficial. This observation aligns with a key characteristic of GPICL, where learning from context takes precedence over zero-shot generalization. 3. For greater complexity ($n>6$), the ICL horizon clearly extends far beyond $T=4k$, and the improvement in perplexity exhibits polynomial scaling law (depicted as a linear relationship on a semi-logarithmic axis) with increasing context length. However, it remains difficult to assert from current experiments whether, given sufficiently long contexts, models of varying scales will converge to similar asymptotic performances. Nevertheless, we can confidently conclude that for $n<8$, a model with 10 million parameters has been adequate. 4. For the PG-19 test set, we were surprised to find that the tiny-sized model achieves comparable asymptotic performance to the other two models. We must consider that the complexity of natural language far exceeds that of any conceivable $n$-gram language. Furthermore, the improvement does not exhibit a linear trend on a semi-logarithmic scale. It is reasonable to hypothesize that the GPICL is only capturing certain short-term and high-frequency patterns in natural language (such as frequently appeared words).
@@ 3.2 Baseline Solution
The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters). A overview of the model architecture is in Figure 7. To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning. To avoid the growing “compounding errors” (Ross & Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011). We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details). Each episode has $2,048$ steps, which gives $26$M frames in total.
@@ 3.4 Static Evaluation
The static evaluation was conducted using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames). To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$). In Figure 8, We compare the performance of the small-sized and standard-sized causal models, with several noteworthy observations: First, the future prediction task shows no clear dependency on long contexts (the learning horizon is below 100 steps), whereas the performance of action prediction continually improves with longer contexts. This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL. This is unsurprising, as world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning. Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model. This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
3.5 Interactive Evaluation and Case Study
@@
We further investigate the online performance of the model by directly interacting with the environment using the learned model. The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$. A softmax policy, based on the predicted distribution of actions, is implemented. We plot the mean and variance of the average accumulated rewards over time (t). For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, and agents with different short-term memory to long-term memory transition probabilities $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$ (Figure 9). The results indicate that our model significantly outperforms random exploration and rule-based agents with low long-term memory capabilities ($p(\text{STM}\rightarrow\text{LTM})=5\%$). However, there is a noticeable gap compared to those with more effective long-term memories ($p(\text{STM}\rightarrow\text{LTM})=25\%,100\%$). Moreover, the rate of improvement is not significant compared to those with effective long-term memories, where we would expect to see an accelerating accumulation of rewards. We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
Figure 10: Randomly selected clips consisting of four continuous frames from the evaluated tasks (generated by a standard-size causal model) are displayed in sequence from top to bottom. The images from left to right: initial observations, observations reconstructed by the VAE, ground truth observations for the next frame, and predicted next frame.
In Figure 10, we present the performance of next-frame predictions for randomly selected clips of four frames generated by the causal decision model. The causal modeling successfully predicts most of the maze layouts in the next frame (including textures on the walls), but occasionally misses distant details. Those distant details contribute less to losses and are more dependent on long-term observations, making them more challenging to acquire. It is also worth noting that the failure to predict the next command (as seen in the second case) is frequent but expected, since the commands are issued randomly.
@@ 4.2 Benchmarks for Meta-Learning and ICL
A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016). Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyper-parameters, such as target positions and robot configurations. Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations.
A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations. The setup of these tasks can appear forced and unrealistic. For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020). In practical settings, such information like targets and directions would typically be explicitly indicated through instructions or visual cues, making the tasks ostensibly complex yet fundamentally simplistic. Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations. Our benchmarks, in contrast, align more closely with real-world challenges and include tasks such as indoor navigation, planning, Simultaneous Localization and Mapping (SLAM), and language modeling. These are designed to mirror actual scenarios where long-term interaction and complex decision-making are critical.
@@ 5 Conclusions
This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (especially for long-term dependency modeling) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.

### v2

@@ Abstract
In-context learning (ICL) capabilities are becoming increasingly appealing for building general intelligence due to their sample efficiency and independence from artificial optimization skills. To enhance generalization, biological neural systems primarily inherit learning capabilities and subsequently refine their memory, acquiring diverse skills and knowledge through extensive lifelong experiences. This process gives rise to the concept of general-purpose in-context learning (GPICL). Compared to standard ICL, GPICL addresses a broader range of tasks, extends learning horizons, and starts at a lower zero-shot baseline. We introduce two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark includes a vast number of tasks characterized by significant task variance and minimal transferable knowledge among tasks, facilitating lifelong in-context learning through continuous generation and interaction. These features pose significant challenges for models that rely on context or interactions to improve their proficiency, including language models, decision models, and world models. Our experiments reveal that parameter scale alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds to: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: source code available at https://github.com/FutureAGI/L3C
@@ 1 Introduction
The community has long recognized a significant gap between artificial intelligence and biological intelligence, particularly regarding innate abilities and the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). For example, the initial limited capabilities of mammal infants stand in stark contrast to their skills in lifelong learning and adaptation. This disparity is further highlighted by the phenomenon known as the genomics bottleneck. Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch and Schmidhuber, 2021; Wang et al., 2022). A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 2. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
@@
Why GPICL? The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models. First, inspired by the genomics bottleneck and building on previous work (Kirsch and Schmidhuber, 2021), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills. Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model. In this context, GPICL could offer an ideal solution: since the meta-training stage can result in the loss of much task-specific knowledge and skills but preserves the core learning functionality, it helps prevent the meta-model from being overly influenced by non-factual information and bias in those tasks. Third, in practice, a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, making it challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
Building on those points, there can be an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans. A visual illustration of this proposed pipeline with existing methodologies is depicted in Figure 2.
@@
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL. One crucial conclusion from these limited experiments may be that parameter scale alone is not crucial for ICL or GPICL. This finding aligns with previous research indicating that the size of contexts and memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022). This insight may highlight alternative approaches beyond merely scaling up models within the current AI community.
@@ 2.3.1 Towards Meta-Language Model
We test the ability of generation of the meta-language model on two rudimentary real-world language tasks: English vocabulary learning and basic mathematical operations (addition and subtraction), which are typical for primary school education. For the English vocabulary task, the vocabulary consists of alphabet characters and common punctuation marks; for the math task, it includes digits and mathematical operators. The input contexts for testing included 40 English words and simple arithmetic operations below $5$. To discourage direct replication, we intentionally introduced errors into $15\%$ of the words or equations. We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts. The results, depicted in Fig. 4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings. Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
@@ 2.3.2 Scaling Laws for GPICL
We have conducted further research into how the scale of parameters affects GPICL capability. As illustrated in Fig. 5, our findings reveal notable trends: 1. Language complexity ($n$) influences the relative performance of different-sized models and affects the dependence on context length. In simpler tasks ($n<4$), performance rapidly converges to lower bounds asymptotically ($t\sim T=4k$. Here, with a slight abuse of notation, $T$ does not necessarily represent the ICL horizon, but rather denotes the maximum context length used in our experiments.), and the performances of various models become indistinguishable, despite starting differently. 2. It should be noted that for tasks as straightforward as $n=2$, where a 300K model achieves comparable asymptotic performance to a 150M model, we observe that larger models cannot attain higher performance without sufficient context. It is reasonable to infer that increasing model complexity further will not be beneficial. This observation aligns with a key characteristic of GPICL, where learning from context takes precedence over zero-shot generalization. 3. For greater complexity ($n>6$), the ICL horizon clearly extends far beyond $T=4k$, and the improvement in perplexity exhibits polynomial scaling law (depicted as a linear relationship on a semi-logarithmic axis) with increasing context length. However, it remains difficult to assert from current experiments whether, given sufficiently long contexts, models of varying scales will converge to similar asymptotic performances. Nevertheless, we can confidently conclude that for $n<8$, a model with 10 million parameters has been adequate. 4. For the PG-19 test set, we were surprised to find that the tiny-sized model achieves comparable asymptotic performance to the other two models. We must consider that the complexity of natural language far exceeds that of any conceivable $n$-gram language. Furthermore, the improvement does not exhibit a linear trend on a semi-logarithmic scale. It is reasonable to hypothesize that the GPICL is only capturing certain short-term and high-frequency patterns in natural language (such as frequently appeared words).
@@ 3.2 Baseline Solution
The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters). A overview of the model architecture is in Fig. 7. To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning. To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011). We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details). Each episode has $2,048$ steps, which gives $26$M frames in total.
@@ 3.3 Baseline Performances
Static Evaluation. We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames). To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$). In Figure 8, We compare the performance of the small-sized and standard-sized causal models, with several noteworthy observations: First, the future prediction task shows no clear dependency on long contexts (the learning horizon is below 100 steps), whereas the performance of action prediction continually improves with longer contexts. This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL. This is unsurprising, since single-step world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning. Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model. This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
@@
Interactive Evaluation. It is more beneficial to investigate the online performance of the policy model and world model in causal modeling by directly interacting with the environment.
We first directly evaluate the model-free policy by interacting with a set of evaluation tasks. The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$. A softmax policy, based on the predicted distribution of actions, is implemented. We plot the mean and variance of the average accumulated rewards over time (t). For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, and agents with different short-term memory to long-term memory transition probabilities $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$ (Figure 9). The results indicate that our model significantly outperforms random exploration and rule-based agents with low long-term memory capabilities ($p(\text{STM}\rightarrow\text{LTM})=5\%$). However, there is a noticeable gap compared to those with more effective long-term memories ($p(\text{STM}\rightarrow\text{LTM})=25\%,100\%$). Moreover, the rate of improvement is not significant compared to those with effective long-term memories, where we would expect to see an accelerating accumulation of rewards. We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
Figure 10: Randomly selected clips depicting the future 8 steps using the world model of causal modeling are presented. For each clip, the upper section shows the ground truth observations from the current step to 8 steps ahead (from left to right), while the bottom section shows the imagined future up to 8 steps ahead starting from the current observation. The actions taken are also labeled with text in the image. The future imagination is generally accurate for the first frame but deteriorates over time due to error propagation, illustrating the challenge of causal world modeling.
We then provide a qualitative evaluation of the performance of world models in Figure 10. Unlike $\mathcal{L}_{wm}$, which can only represent the precision of predicting one-step transitions, using the world model to predict the distant future is far more challenging. We observe that although our world model can predict the nearest future, its accuracy quickly deteriorates over time. This can be attributed to the well-known problem of horizon-square compounding errors. Typical improvements include adding noise to the inputs (Hu et al., 2022) and multi-step prediction (Asadi et al., 2019). However, compared to many previous problems, Maze World poses a higher challenge to world modeling in two key aspects: First, the variance in observations is more severe (the observation of two consecutive frames can be very different). Second, the variance of the tasks requires the world model to perform in-context learning to improve, as each maze is completely different and not transferrable.
@@ 4.2 Benchmarks for Meta-Learning and ICL
A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016). Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyper-parameters, such as target positions and robot configurations. Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations. Mazes are also frequently used to benchmark meta-reinforcement learning (Duan et al., 2016; Mishra et al., 2018), but the scale and diversity of those tasks are not adequate enough to benchmark GPICL.
A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations. For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020). In practical settings, such information like targets and directions would typically be explicitly indicated through instructions or visual cues, making the tasks ostensibly complex yet fundamentally simplistic. Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations. Our benchmarks, in contrast, align more closely with real-world challenges. For instance, the Maze World “navigation” tasks can mirror the adaptive learning process of a household robot adapting to a new indoor scenario, where it needs to simultaneously solve localization, mapping, and planning. Another group of Maze World “survival” tasks simulate an agent exploring a new environment to learn about helpful and harmful elements.
@@ 5 Conclusions
This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.